• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name pking

    Log entry time 00:04:13 on April25,2001

    Entry number 61303

    keyword=Swing shift summary.

    Hyper-white scan run summary:

    16:55 The 15cm LH2 is in position. Request 90 uA of beam using the 2x2
    raster, centered at X=+1.0 , Y=0.0.
    17:15 Run 3879 is a test of the prescale factors;the right dipole had not settled
    to 1.065 GeV/c yet.
    17:35 The collimators for both arms have been set to "open". The right dipole
    has still not yet settled to 1.065 GeV/c.
    17:55 The right spectrometer has reached 1.065 GeV/c.
    Begin run 3880: Momentum setting #1, open collimators, LH2 target.
    18:14 MCC lost an IOC; beam will be down for a little while.
    We stop run 3880, and start changing the collimators.
    18:24 Both collimators are set at 6 msr; beam has not been restored.
    19:00 Beam is back; IOC problem at MCC fixed.
    Begin run 3881: Momentum setting #1, 6 msr collimators, LH2 target.
    19:25 End run 3881. Ramp the spectrometers down to 1.0437 GeV/c.
    19:35 Spectrometers are stabilized at 1.0437 GeV/c.
    Begin run 3882: Momentum setting #2, 6 msr collimators, LH2 target.
    20:00 End run 3882.
    Move the collimators to the "open" position.
    20:08 Collimators are "open".
    Begin run 3883: Momentum setting #2, open collimation, LH2 target.
    20:28 End run 3883.
    Set spectrometers to 1.0228 GeV/c.
    20:38 Spectrometers are stabilized at 1.0228 GeV/c.
    Begin run 3884: Momentum setting #3, open collimation, LH2 target.
    20:58 End run 3884.
    Move to the 9-foil carbon target.
    21:02 Request beam with the raster turned off.
    21:04 Begin run 3885: Momentum setting #3, open collimation, optics target,
    raster turned off.
    21:10 End run 3885.
    Change the collimation to 6 msr.
    21:17 Collimators are set to 6 msr.
    Begin run 3886: Momentum setting #3, 6 msr collimation, optics target,
    raster is turned off.
    21:23 There were ROC communication errors during run 3886. All of the ROCs
    seemed to be alive, though.
    Begin run 3887 with the same conditions.
    21:29 End run 3887.
    Move to the 15 cm LH2 target.
    21:32 The LH2 target is in position. Request beam with the 2x2 raster pattern.
    21:33 Beam is on.
    Begin run 3888: Momentum setting #3, 6 msr collimation, LH2 target.
    21:44 The computer deadtime is going up to around 60%.
    End run 3888.
    Begin run 3889: Momentum setting #3, 6 msr collimation, LH2 target.
    The deadtime is back to about 18% now.
    22:00 End run 3889.
    Ramp spectrometers to 1.0023 GeV/c.
    22:08 Spectrometers have stabilized at 1.0023 GeV/c.
    Begin run 3890: Momentum setting #4, 6 msr collimation, LH2 target.
    22:31 End run 3890.
    Change collimation to the "open" position.
    22:37 Collimators are in the "open" position on both spectrometers.
    Begin run 3891: Momentum setting #4, open collimation, LH2 target.
    22:54 End run 3891.
    Change the spectrometer momenta to 0.9823 GeV/c, and change to the
    9-foil optics target.
    22:58 Optics target is in position, request beam with the raster off; the spectrometers
    have not quite stabilized yet.
    23:04 The spectrometers have stabilized at 0.9823 GeV/c.
    Begin run 3892: Momentum setting #5, open collimation, optics target,
    raster turned off.
    23:09 End run 3892.
    Change spectrometer collimation to 6 msr.
    23:15 Spectrometer collimation is 6 msr on both arms.
    Begin run 3893: Momentum setting #5, 6 msr collimation, optics target,
    raster is turned off.
    23:20 Move to the LH2 target.
    23:25 The 15 cm LH2 target is in position. Request beam with the 2x2 raster
    pattern.
    Begin run 3894: Momentum setting #5, 6 msr collimation, LH2 target.
    23:50 End run 3894.
    Change spectrometer collimation to "open".
    23:55 The collimators are in "open" position.
    We are ready to take the last data run, but there is suddenly a problem in the
    injector. MCC cannot deliver beam to us right now; they have called the
    injector-on-call.

    The last data run would be the open collimation run on LH2 at this
    spectrometer setting of 0.9823 GeV/c.